The shearing stress of a certain steel alloy is 2.5 x 10^8 Pa. A 5.0 mm diameter cylindrical bolt of this alloy is used to fasten a bracket to a wall. What is the minimum load the bracket can support without shearing off the bolt?
1
Expert's answer
2020-02-11T10:11:43-0500
"F=PA=0.25\pi d^2P"

"F=0.25\pi (0.005)^2(2.5 \cdot 10^8)=4900\ N=4.9\ kN"